# Multi-label Scene Classification for Autonomous Vehicles: Acquiring and Accumulating Knowledge from Diverse Datasets

### v2

@@ IV Methodology
Multi-label driving scene classification involves assigning multiple non-exclusive class labels to each input scene image $\boldsymbol{x}$ in the domain =$\mathcal{I}=\{\boldsymbol{x}|\boldsymbol{x}\in\mathbb{R}^{3\times H\times W}\}$, a set of color images with three color channels, a height of $H$ pixels, and a width of $W$ pixels. $\boldsymbol{x}$ can be classified based on $M$ different scene attributes, $Y_{m}$, for $m=1,\dots,M$. Each attribute $Y_{m}$ is a categorical variable with its support $\mathcal{U}_{m}$ comprising the exclusive classes for this attribute. A multi-label classification model $\mathcal{C}$ approximates the mapping, $\mathcal{I}\rightarrow\mathcal{U}_{1}\times\dots\times\mathcal{U}_{M}$, which estimates the class labels for each scene image $\boldsymbol{x}$:
$[\widehat{\boldsymbol{y}}_{1},\dots,\widehat{\boldsymbol{y}}_{M}]=\mathcal{C}(\boldsymbol{x};\boldsymbol{\Theta}),$ (2)
where $\widehat{\boldsymbol{y}}_{m}\in\mathbb{R}^{|\mathcal{U}_{m}|}$ is the prediction in probability for the one-hot encoded true label $\boldsymbol{y}_{m}\in\mathbb{B}^{|\mathcal{U}_{m}|}$ of $\boldsymbol{x}$, and $\boldsymbol{\Theta}$ represents the learnable parameters of the classification model $\mathcal{C}$. To build such a model, this section proposes a new deep learning method named KAA-CAL. First, KAA learns a model $\widetilde{\mathcal{C}}=\mathcal{C}(\cdot,\widetilde{\boldsymbol{\Theta}})$ that maximizes the classification accuracy in isolated per-attribute identification. Following this, CAL further adapts $\widetilde{\mathcal{C}}$ to become $\mathcal{C}^{\ast}=\mathcal{C}(\cdot,\boldsymbol{\Theta}^{\ast})$ that is proficient in simultaneous all-attribute identification.
Figure 2: Overview of the proposed Knowledge Acquisition and Accumulation learning system (KAA)
@@ IV-A Knowledge Acquisition and Accumulation (KAA)
KAqN’s deep encoder is trained to attain the knowledge for performing $M$ downstream tasks, indexed by $m$, each focused on classifying the input driving scene image according to one attribute. The overall approach of KAqN is a sequential and cyclical learning process, as outlined in Algorithm 1. A training epoch, indexed by $t$, is one learning cycle. During each learning cycle, KAqN goes through $M$ learning iterations to sequentially learn on these classification tasks, utilizing a dedicated single-label dataset for each task. Mathematically speaking, the current learning iteration $i$ corresponds to a specific learning cycle $t$ and learning task $m$:
@@
Conversely, the training on task $m$ within cycle $t$ is conducted in learning iteration $i$:
$i=(t-1)M+m.$ (4)
@@
       KAqN acquires new knowledge sequentially:
@@
      $\boldsymbol{\theta}(Mt)$ is consolidated with $\boldsymbol{\theta}^{{}^{\prime}}(t-1)$:
@@
       Termination of the learning process:
@@
OUTPUT: $\widetilde{\boldsymbol{\Theta}}\leftarrow\{\boldsymbol{\theta}^{{}^{\prime}}(\tilde{t}),\boldsymbol{\phi}_{1}(\tilde{t}),\dots,\boldsymbol{\phi}_{M}(\tilde{t})\}$
@@
IV-A1 Knowledge Acquisition
KAqN utilizes the backbone of a Swin-B [27], a vision transformer pretrained on ImageNet, as the deep feature encoder. The encoder embeds each input image as a feature map for the downstream classification tasks. Specifically, Swin-B processes an input scene image with a resolution of $244\times 244$ by first partitioning it into non-overlapping $4\times 4$ patches. It then sequentially encodes these patches into a feature map through four stages. Each stage applies either a linear embedding ($\operatorname{LE}$) or a patch merging ($\operatorname{PM}$) layer, followed by multiple Swin Transformer ($\operatorname{ST}$) blocks:
@@ IV-A1 Knowledge Acquisition
Here, $\boldsymbol{\theta}^{\mathcal{E}}$ designates all the learnable parameters of KAqN’s deep encoder in Eq. (5). The extracted feature maps have the following dimensions: $\boldsymbol{f}_{\textsc{i}}\in\mathbb{R}^{128\times 56\times 56}$, $\boldsymbol{f}_{\textsc{ii}}\in\mathbb{R}^{256\times 28\times 28}$, $\boldsymbol{f}_{\textsc{iii}}\in\mathbb{R}^{512\times 14\times 14}$, and $\boldsymbol{f}_{\textsc{iv}}\in\mathbb{R}^{1024\times 7\times 7}$. That is, the encoder in Eq. (5) progressively reduces the spatial resolution of the input scene image while increasing the depth (number of channels) of the feature maps.
Then, four projectors independently embed the individual feature maps derived from each stage as feature vectors,
@@
The fully-connected (FC) layer of the classifier reduces the dimension of feature embedding $\boldsymbol{d}_{\textsc{iv}}$ to match the class size of attribute $m$, and the softmax operator ($\operatorname{SM}$) normalizes the output as the probability distribution over classes. $\boldsymbol{\phi}_{m}$ represents the learnable parameters of classifier $m$ in Eq. (7).
@@
where $\cdot$ denotes the inner product of the one-hot encoded ground truth class label, $\boldsymbol{y}_{m}$, and the predicted probability distribution over classes, $\widehat{\boldsymbol{y}}_{m}$. The prediction loss in Eq. (8) is part of the total loss (to be introduced in Eq. (15)) for guiding the modeling training.
Within a learning cycle $t$, KAqN learns classification tasks one at a time, using a dedicated single-label dataset for each task. In completing a learning iteration described in Eqs. (5, 6, 7, 15) , KAqN’s parameter values are updated:
[…]
@@
[…]
@@
IV-A2 Knowledge Accumulation
Knowledge acquired by KAqN in the cyclical training process is accumulated in KAcN. Fig. 2 shows that KAcN’s network architecture is identical to KAqN. The feature map extracted in each stage of the KAcN’s encoder is $\boldsymbol{f}^{{}^{\prime}}_{s}$, which is further embedded as a feature vector $\boldsymbol{d}^{{}^{\prime}}_{s}$, for $s=\textsc{i},\dots,\textsc{iv}$. For simplicity, $\boldsymbol{\theta}^{{}^{\prime}}$ designates all the learnable parameters of KAcN. The identical network architecture of KAcN and KAqN greatly simplifies the process of knowledge accumulation delineated as below.
@@ IV-A2 Knowledge Accumulation
where $T$ is the maximum learning cycles for KAA. Eq. (12) indicates that $\alpha(t)$ is a monotonically increasing function. That is, in the early phases of learning, KAA leans toward plasticity, allowing itself to rapidly absorb new knowledge. In the later phases, it prioritizes stability and thus becomes more conservative in absorbing new knowledge to prevent the disruption of the established cognitive capability.
IV-A3 Knowledge Retention
The cyclical nature of training can cause catastrophic forgetting, thus impeding continual learning and improvement. This issue can be effectively addressed by setting the accumulated knowledge as the reference for the subsequent knowledge acquisition. KAA achieves this through feature-based KD, which penalizes stage-wise discrepancies between the feature embedding extracted by KAqN, $\boldsymbol{d}_{s}(i)$, in each learning iteration and the corresponding feature embedding derived by KAcN, $\boldsymbol{d}^{\prime}_{s}(t-1)$. The penalty is calculated as the mean squared error (MSE):
@@ IV-A3 Knowledge Retention
for stages $s=\textsc{I},\dots,\textsc{IV}$. This particular KD design fully leveraging the advantages of network depth has been shown to be effective [19].
@@
where $\lambda_{s}$ is the coefficient for stage $s$. In this study, $\lambda_{s}$ is 1 for all stages, meaning equal weights are assigned across stages.
IV-A4 Loss Function as Learning Guidance
KAA uses the consistency loss calculated in Eq. (14) as a regularization term for training KAqN to ensure the new knowledge acquisition is based on already attained knowledge. Therefore, the per-sample loss function in learning iteration $i$ can be formulated as the weighted sum of the prediction loss and the consistency loss:
@@ IV-A4 Loss Function as Learning Guidance
where $\beta(i)$, a real-valued parameter in the range $[0,1]$, is the performance-based acquisition-retention indicator (PARI) for balancing the acquisition of new knowledge and the retention of already attained knowledge. $\beta(i)$ is designed to be adjusted dynamically with respect to the learning needs. Lower performance anticipated for KAqN on the task in learning corresponds to a greater need for continual learning and improvement on this task. Therefore, $\beta(i)$ is designed as a monotonically decreasing function of the estimated task performance, $\widehat{p}(i)$:
@@
where $\psi>1$ is a real-valued parameter defining the shape of the curve. In this study, $\psi$ is set to be 4.
In Eq. (16), $\widehat{p}(i)$, the estimated classification accuracy of KAqN in learning iteration $i$, is obtained based on an EMA process:
@@
where $\widehat{p}(i-M)$ is the estimated accuracy of KAqN for the last learning cycle and $p(i-M)$ is the accuracy that KAqN actually achieved. When a prior measurement or estimation is not available, a default value, $\widehat{p}(0)$, representing KAqN’s initial (low) performance, is used as a substitute. The parameter $\omega$ is set to be $0.9$ in this study, meaning that the performance estimation puts more weight on the long-term trend than the most recent observation.
@@ IV-B Consistency-based Active Learning (CAL)
KAA learns to perform multi-label scene classification through monotask learning across different single-label datasets. While simplifying data collection and reducing learning complexity, this approach encounters a domain shift issue. Precisely, the class label distributions of single-label datasets represent the marginal probabilities of individual attributes, different from their joint probabilities. A similar domain shift occurs in the extracted features. Consequently, the resulting classification model can make mistakes when it is tested on samples drawn from the joint distribution of driving scenes.
@@
This paper proposes CAL to tackle the aforementioned issue, which seeks to accomplish multi-label classification in real-world scenarios in a cost-effective manner. Eventually, the classification model obtained by KAA via monotask learning will adapt to multitask learning, thereby achieving satisfactory performance for multi-label classification.
Fig. 3 illustrates the overall approach of the proposed CAL. A multi-label test sample, $\mathcal{D}^{T}$, is prepared, representing the driving scene distribution in the joint space of scene attributes. As outlined in Fig. 3, a stratified sampling strategy ($\operatorname{Stf}$) is implemented, which randomly samples $\kappa$ test images per class from DSI’s test sets:
$\mathcal{D}^{T}=\cup_{m=1}^{M}\operatorname{Stf}(\mathcal{D}_{m}^{Tst};\mathcal{U}_{m},\kappa),$ (18)
which is then fully annotated with respect to all $M$ attributes.
Then, a labeled training set $\mathcal{D}^{L}$ needs to be obtained to adapt the classification model for multitask learning over iterations. Although the ground truth labels for the training set $\mathcal{D}_{m}^{Trn}$ were seen by task $m$, other tasks have never seen these labels. Therefore, the union of DSI’s training sets forms an “unlabeled” pool,
$\mathcal{D}^{U}=\cup_{m=1}^{M}\mathcal{D}_{m}^{Trn},$ (19)
@@
CAL, as outlined in Fig. 3 and Algorithm 2, searches the unlabeled pool $D^{U}$ for data points that are similar to those in the test set $D^{T}$. A consistency function, $\operatorname{con}(v,u)$, is defined to quantify the similarity between any image in the test set $v\in\mathcal{D}^{T}$ and any one in the unlabeled pool $u\in\mathcal{D}^{U}$:
$\operatorname{con}(v,u)=-\|v,u\|_{\boldsymbol{d}^{{}^{\prime}}_{\textsc{iv}},2},$ (20)
where $\|\cdot,\cdot\|_{\boldsymbol{d}^{{}^{\prime}}_{\textsc{iv}},2}$ measures the Euclidean distance between two data points based on their final-stage feature embeddings. Thereby, for any image in the test set $v\in\mathcal{D}^{T}$, its nearest neighbor (i.e., the corresponding image in the unlabeled pool which is most similar to it) is found:
@@
[…]
where $\boldsymbol{\Theta}^{{}^{\prime}}(0)=\widetilde{\boldsymbol{\Theta}}$. This process can go for multiple iterations until satisfaction, resulting in the optimized classification model $\mathcal{C}^{\ast}$ parameterized with $\boldsymbol{\Theta}^{\ast}$.
INPUT:
$\mathcal{D}^{U}$: Unlabeled pool
$\mathcal{D}^{T}$: Test dataset
$b_{j}$: Budget for adaptation iteration $j$
$N_{\text{CAL}}$: Maximum iterations
INITIALIZATION:
$\mathcal{D}^{L}=\emptyset$: initial labeled training set
CONSISTENCY ACTIVE LEARNING:
for $j=1,\dots,N_{\text{CAL}}$ do
      add $\mathcal{A}(j)$ to the labeled dataset:
       $\mathcal{D}^{L}\leftarrow\mathcal{D}^{L}\,\cup\,\mathcal{A}(j)$;
      remove $\mathcal{A}(j)$ from unlabeled dataset:
       $\mathcal{D}^{U}\leftarrow\mathcal{D}^{U}\setminus\mathcal{A}(j)$;
      train the classification model for multitasking:
       if stopping criterion is satisfied then
            $j^{*}=j$.
             break
       end if
end for
Output: $\boldsymbol{\Theta}^{\ast}=\boldsymbol{\Theta}^{{}^{\prime}}(j^{\ast})$
Algorithm 2 CAL Algorithm

### v3

@@ IV Methodology
Multi-label driving scene classification involves assigning multiple non-exclusive class labels to each input scene image $\boldsymbol{x}$ in the domain $\mathcal{I}=\{\boldsymbol{x}|\boldsymbol{x}\in\mathbb{R}^{3\times H\times W}\}$, comprising color images with three color channels, a height of $H$ pixels, and a width of $W$ pixels. $\boldsymbol{x}$ can be classified based on $M$ different scene attributes, $Y_{m}$, for $m=1,\dots,M$. Each attribute $Y_{m}$ is a categorical variable with its support $\mathcal{U}_{m}$ comprising the exclusive classes for this attribute. A multi-label classification model $\mathcal{C}$ approximates the mapping, $\mathcal{I}\rightarrow\mathcal{U}_{1}\times\dots\times\mathcal{U}_{M}$, which estimates the class labels for each scene image $\boldsymbol{x}$:
$\widehat{\boldsymbol{y}}_{1},\dots,\widehat{\boldsymbol{y}}_{M}]=\mathcal{C}(\boldsymbol{x};\boldsymbol{\Theta}),$ (2)
[…]
Figure 3: Overview of the proposed Knowledge Acquisition and Accumulation (KAA)
learning system
@@ IV-A1 Overview of the Learning Process
KAqN’s deep encoder is trained to attain the knowledge for performing $M$ downstream tasks, indexed by $m$, each focused on classifying the input driving scene image according to one attribute. The overall approach of KAqN is a sequential and cyclical learning process, as outlined in Algorithm 1. A training epoch, indexed by $t$, is one learning cycle. During each learning cycle, KAqN goes through $M$ learning iterations to sequentially learn on these classification tasks, utilizing a dedicated single-label dataset for each task. Therefore, the learning iteration $i$ corresponds to a specific learning cycle $t$ and learning task $m$:
@@
  1exKAqN acquires new knowledge sequentially:
@@
[…]
@@
  1exTermination of the learning process:
@@
[…]
@@
IV-A2 Knowledge Acquisition
KAqN utilizes the backbone of Swin-B [25], a vision transformer pretrained on ImageNet, as the deep feature encoder. The encoder embeds each input image as a feature map for the downstream classification tasks. Specifically, Swin-B processes an input scene image with a resolution of $244\times 244$ by first partitioning it into non-overlapping $4\times 4$ patches. It then sequentially encodes these patches into a feature map through four stages, progressively reducing the spatial resolution of the input scene image while increasing the depth (number of channels) of the feature maps. Each stage applies either a linear embedding ($\operatorname{LE}$) or a patch merging ($\operatorname{PM}$) layer, followed by multiple Swin Transformer ($\operatorname{ST}$) blocks:
@@ IV-A2 Knowledge Acquisition
Here, $\boldsymbol{\theta}^{\mathcal{E}}$ designates all the learnable parameters of KAqN’s deep encoder. The extracted feature maps have the following dimensions: $\boldsymbol{f}_{\textsc{i}}\in\mathbb{R}^{128\times 56\times 56}$, $\boldsymbol{f}_{\textsc{ii}}\in\mathbb{R}^{256\times 28\times 28}$, $\boldsymbol{f}_{\textsc{iii}}\in\mathbb{R}^{512\times 14\times 14}$, and $\boldsymbol{f}_{\textsc{iv}}\in\mathbb{R}^{1024\times 7\times 7}$.
Then, four projectors respectively embed the individual feature maps derived from each stage as feature vectors,
@@
The fully-connected (FC) layer of the classifier reduces the dimension of feature embedding $\boldsymbol{d}_{\textsc{iv}}$, and the softmax operator ($\operatorname{SM}$) normalizes the output as the probability distribution over classes. $\boldsymbol{\phi}_{m}$ in Eq. (6) represents the learnable parameters of classifier $m$.
@@
where $\cdot$ denotes the inner product of the one-hot encoded ground truth class label, $\boldsymbol{y}_{m}$, and the predicted probability distribution over classes, $\widehat{\boldsymbol{y}}_{m}$. The prediction loss in Eq. (7) is part of the total loss for guiding the modeling training.
Within a learning cycle $t$, KAqN learns classification tasks one at a time, using a dedicated single-label dataset for each task. In completing a learning iteration described in Eqs. (4, 5, 6, 14) , KAqN’s parameter values are updated:
$\boldsymbol{\theta}(i)\xleftarrow{\text{monotask learning}}\boldsymbol{\theta}(i-1),$ (8)
@@
$\boldsymbol{\phi}_{m}(t)\xleftarrow{\text{monotask learning}}\boldsymbol{\phi}_{m}(t-1).$ (9)
@@
IV-A3 Knowledge Accumulation
Knowledge acquired by KAqN in the cyclical training process accumulates in KAcN. Fig. 3 shows that KAcN’s network architecture is identical to KAqN. The feature map extracted in each stage of the KAcN’s encoder is $\boldsymbol{f}^{{}^{\prime}}_{s}$, which is further embedded as a feature vector $\boldsymbol{d}^{{}^{\prime}}_{s}$, for $s=\textsc{i},\dots,\textsc{iv}$. For simplicity, $\boldsymbol{\theta}^{{}^{\prime}}$ designates all the learnable parameters of KAcN. The identical network architecture of KAcN and KAqN greatly simplifies the process of knowledge accumulation delineated as below.
@@ IV-A3 Knowledge Accumulation
where $T$ is the maximum learning cycles for KAA. Eq. (11) indicates that $\alpha(t)$ is a monotonically increasing function. That is, in the early phases of learning, KAA leans toward plasticity, allowing itself to rapidly absorb new knowledge. In the later phases, it prioritizes stability to prevent the disruption of the established cognitive capability.
IV-A4 Knowledge Retention
The cyclical nature of training can cause catastrophic forgetting, thus impeding continual learning and improvement. This issue can be effectively addressed by setting the accumulated knowledge as the reference for the subsequent knowledge acquisition. KAA achieves this through feature-based KD, which penalizes large stage-wise discrepancies between the feature embedding extracted by KAqN, $\boldsymbol{d}_{s}(i)$, in each learning iteration and the corresponding feature embedding derived by KAcN, $\boldsymbol{d}^{\prime}_{s}(t-1)$. The penalty is calculated as the mean squared error (MSE):
@@ IV-A4 Knowledge Retention
for stages $s=\textsc{I},\dots,\textsc{IV}$. This particular KD design fully leveraging the advantages of deep networks, which has been shown to be effective [18].
@@
where $\lambda_{s}$ is the coefficient for stage $s$, which is set to be one for all stages in this study.
IV-A5 Loss Function as the Learning Guidance
The consistency loss in Eq. (13) is a regularization term for training KAqN, ensuring that the new knowledge acquisition is based on already attained knowledge. Therefore, the per-sample loss function in learning iteration $i$ can be formulated as the weighted sum of the prediction loss and the consistency loss:
@@ IV-A5 Loss Function as the Learning Guidance
where $\beta(i)$, a real-valued parameter in the range $[0,1]$, is the performance-based acquisition-retention indicator (PARI) for balancing the acquisition of new knowledge and the retention of already attained knowledge. $\beta(i)$ is adjusted dynamically with respect to learning needs. Lower performance anticipated for KAqN on the task in learning corresponds to a greater need for continual learning and improvement on this task. Therefore, $\beta(i)$ is designed as a monotonically decreasing function of the estimated task performance, $\widehat{p}(i)$:
@@
where $\psi>1$ is a real-valued parameter defining the shape of the curve, which is set to be 4 in this study.
$\widehat{p}(i)$ in Eq. (15) estimates classification accuracy of KAqN in learning iteration $i$, obtained based on an EMA process:
@@
where $\widehat{p}(i-M)$ is the estimated accuracy of KAqN for the last learning cycle and p(iM)p(i-M) is the accuracy that KAqN actually achieved. The EMA process ensures that the performance estimation is based on the entire record from past training cycles. The parameter $\omega$ is set to be $0.9$ in this study, meaning that the performance estimation puts more weight on the long-term trend than the most recent observation.
@@ IV-B Consistency-based Active Learning (CAL)
KAA learns the knowledge for multi-label scene classification through monotask learning across different single-label datasets. While simplifying data collection and reducing learning complexity, this approach confronts a domain shift issue. That is, the class label distributions of single-label datasets represent the marginal probability distributions of individual attributes, different from their joint distribution. The same issue occurs in the extracted features. Therefore, the foundation model can make mistakes when it attempts to simultaneously identify a scene using multiple attributes.
@@
INPUT:
$\mathcal{D}^{U}$: Unlabeled pool
$\mathcal{D}^{T}$: Test dataset
$b_{j}$: Budget for adaptation iteration $j$
$N_{\text{CAL}}$: Maximum iterations
1exINITIALIZATION:
$\mathcal{D}^{L}=\emptyset$: initial labeled training set
1exCONSISTENCY ACTIVE LEARNING:
for $j=1,\dots,N_{\text{CAL}}$ do
  add $\mathcal{A}(j)$ to the labeled dataset:
 $\mathcal{D}^{L}\leftarrow\mathcal{D}^{L}\,\cup\,\mathcal{A}(j)$;
  remove $\mathcal{A}(j)$ from unlabeled dataset:
 $\mathcal{D}^{U}\leftarrow\mathcal{D}^{U}\setminus\mathcal{A}(j)$;
  train the classification model for multitasking:
  1exif stopping criterion is satisfied then
    $j^{*}=j$.
    break
   end if
end for
1exOutput: $\boldsymbol{\Theta}^{\ast}=\boldsymbol{\Theta}^{{}^{\prime}}(j^{\ast})$
Algorithm 2 CAL Algorithm
This paper proposes CAL (Algorithm 2) to tackle the aforementioned issue, which seeks to cost-effectively adapt the attained foundation model to real-world multi-attribute scene identification. Fig. 4 illustrates the overall approach of the proposed CAL. A multi-label test sample, $\mathcal{D}^{T}$, representing the driving scene distribution in the joint space of scene attributes, is prepared using a stratified sampling strategy ($\operatorname{Stf}$). It randomly samples $\kappa$ test images per class from DSI’s test sets:
$\mathcal{D}^{T}=\cup_{m=1}^{M}\operatorname{Stf}(\mathcal{D}_{m}^{\text{Tst}};\mathcal{U}_{m},\kappa),$ (17)
which is then fully annotated with respect to all $M$ attributes. The adaptation test set may also be drawn from alternative domains or tasks to which the foundation model has been adapted. The adaptation test set may also be drawn from other domains or tasks to which the foundation model has been adapted.
Then, a training set $\mathcal{D}^{L}$ needs to be labeled for adapting the classification model to multitask learning over iterations. Although the ground truth labels for the training set $\mathcal{D}_{m}^{\text{Trn}}$ were seen by task $m$, other tasks have never seen these labels. Therefore, the union of DSI’s training sets forms an “unlabeled” pool,
$\mathcal{D}^{U}=\cup_{m=1}^{M}\mathcal{D}_{m}^{\text{Trn}},$ (18)
@@
CAL searches the unlabeled pool $D^{U}$ for data points that are similar to those in the test set $D^{T}$. A consistency function, $\operatorname{con}(v,u)$, is defined to quantify the similarity between any image in the test set, $v\in\mathcal{D}^{T}$, and any one in the unlabeled pool, $u\in\mathcal{D}^{U}$:
$\operatorname{con}(v,u)=-\|v,u\|_{\boldsymbol{d}^{{}^{\prime}}_{2,\textsc{iv}}},$ (19)
where $\|\cdot,\cdot\|_{\boldsymbol{d}^{{}^{\prime}}_{2,\textsc{iv}}}$ measures the Euclidean distance between two data points in terms of their final-stage feature embeddings. Thereby, for any image in the test set $v\in\mathcal{D}^{T}$, the corresponding image in the unlabeled pool which is most similar to it is identified:
@@
$\boldsymbol{\Theta}^{{}^{\prime}}(j)\xleftarrow{\text{multitask learning}}\boldsymbol{\Theta}^{{}^{\prime}}(j-1),$ (26)
where $\boldsymbol{\Theta}^{{}^{\prime}}(0)=\widetilde{\boldsymbol{\Theta}}$. This process can go for several iterations until satisfaction, resulting in the optimized classification model $\mathcal{C}^{\ast}$ parameterized with $\boldsymbol{\Theta}^{\ast}$.